OpenStudy (anonymous):

2x^2 - 3x - 2 = 0 (by factoring)

OpenStudy (nikvist):

\[2x^2-3x-2=0\]\[2x^2-4x+x-2=0\]\[2x(x-2)+x-2=0\]\[(x-2)(2x+1)=0\]